### v1

# 208Pb nuclear charge radius revisited: closing the fine-structure-anomaly gap

Abstract
A comprehensive reevaluation of the root-mean-square nuclear charge radius is presented for the doubly magic 208Pb extracted from muonic spectroscopy measurements. By integrating rigorous theoretical quantum electrodynamics calculations, state-of-the-art numerical methods, and a systematic reanalysis of the uncertainties, we reduced the long-standing muonic fine-structure anomaly and improved the goodness of fit by a factor of twenty. The resulting value of 5.5062(5) fm is fairly consistent with the previously reported muonic spectroscopy value, and three standard deviations larger than the commonly used compilation data, which indicates that the current value and its uncertainty could be significantly underestimated. Our study paves a new path for systematic reevaluation of all rms radii based on muonic spectroscopy.
Introduction. — Nuclear root-mean-square (rms) charge radii serve as fundamental benchmarks bridging various fields of physics. They are indispensable input parameters for nuclear-, atomic-, and molecular-physics calculations. Reliable rms radii are crucial for precision tests of quantum electrodynamics (QED) and for the determination of fundamental constants. Additionally, many searches for physics beyond the Standard Model rely on high-precision rms charge radii.
Accurate calculations of nuclear parameters from first principles are particularly challenging for heavy nuclei. Since direct measurements of nuclear charge radii are not possible, indirect extraction with significant theory input is necessary. Muonic atoms are highly sensitive to these nuclear parameters, such that the most accurate method to measure rms charge radii is based on muonic atom spectroscopy [1, 2]. It is also common to combine the values obtained from muonic atoms with electron scattering cross-section data [3] via Barrett moments [4].
Motivated by the success of light muonic atom spectroscopy [5, 6, 7] and new measurements of heavy muonic atoms with microtargets [8], we started theoretical efforts to extract the rms radii of heavy atoms. A literature study revealed the existence of another puzzle, the muonic fine-structure anomaly, discovered about 50 years ago [9, 10, 11, 12] and not mentioned since then: the theoretically predicted spectra of lead, tin, and zirconium exhibit small but systematic deviations from the experimentally measured values, for any value of the assumed nuclear rms charge radius. This discrepancy was attributed to the least understood and most complicated of all contributions: nuclear polarization. However, our rigorous microscopic calculations of the nuclear polarization have shown that it is most likely not solely responsible for the anomaly [13]. This aligns with the poor fit quality addressed above and hints at the uncertainty of the nuclear radius of double-magic 208Pb being significantly underestimated. The same situation applies to other muonic atoms, as the methods for rms charge radius extraction have been the same.
Figure 1: Nuclear rms radius for 208Pb extracted in this work compared to the literature values. The data based on muonic spectroscopy are marked with green (this work (TW) and Bergem [11]), data based on the electron scattering are in red (Mazanek [14] [^2] We actually found this thesis, read and digitalized it. and Dreher [15]), and the combined values are in blue (Angeli [16] and Fricke [17]).
[…]
Theory improvements. — We calculated all individual contributions to the muonic spectrum which contribute $\gtrsim$10of the experimental precision. The leading contribution is the finite nuclear size (FNS) effect, obtained by numerically solving the Dirac equation with a nuclear potential derived from a two-parameter Fermi (2pF) nuclear charge model (see Sup. Mat. for details).
State FNS Ue WK KS $\mu$Ue HadUe SE NP RRecoil Screening
$1s_{1/2}$ 10525.905 67.084 -0.504 0.552 0.238 0.159 -3.345(50) 5.712(603) -2.555(66) 5.555
$2s_{1/2}$ 13581.659 19.336 -0.250 0.149 0.041 0.027 -0.653(6) 1.063(156) -1.317(27) 5.536
$2p_{1/2}$ 14783.784 32.290 -0.356 0.251 0.044 0.030 -0.440(20) 1.923(263) -2.114(4) 5.548
$2p_{3/2}$ 14601.766 29.761 -0.344 0.229 0.033 0.023 -0.745(10) 1.864(256) -2.172(2) 5.547
$3p_{1/2}$ 12129.324 10.759 -0.167 0.081 0.014 0.009 -0.180(15) 0.584(99) -1.003(3) 5.512
$3p_{3/2}$ 12082.436 10.230 -0.163 0.077 0.011 0.008 -0.245(10) 0.619(96) -1.025(0) 5.510
$3d_{3/2}$ 12163.565 10.502 -0.192 0.076 0.001 0.001 -0.020(2) 0.238(56) -1.162(1) 5.525
$3d_{5/2}$ 12121.477 19.848 -0.185 0.071 0.001 $0$0.070(5) 0.037(37) -1.143(0) 5.523
[…]
To the first order in the fine-structure constant $\alpha$, there are two QED corrections: vacuum polarization (VP) and self energy (SE). In contrast to electronic systems, where these two contributions have similar magnitude, VP dominates in muonic atoms. VP can be expressed as a sum of Uehling (Ue) and Wichmann-Kroll (WK) contributions, represented through their respective potentials. The dominant Ue corrections are calculated rigorously for a 2pF model using the method [18]. Since numerical precision is the only limiting factor, we ensured that the error is negligible compared to the experimental uncertainty. For the evaluation of the WK contribution including the FNS effect, two independent methods have been used [19, 20]. Moreover, we have also included the Källén-Sabry (KS) contribution, which accounts for the second-order correction beyond the Uehling term, following the method described in Ref. [21].
In addition to the electron-positron loop, VP can also originate from any other particle-antiparticle (e.g. muon-antimuon) loop. However, these effects are exponentially suppressed by the mass of the virtual particles; thus, going beyond leading order is only necessary for electronic Ue at the current level of accuracy. The muon-antimuon loop gives the muonic Ue ($\mu$Ue) correction, which can be calculated by a straightforward replacement of the virtual particle mass. For the hadronic Ue (HadUe) contribution, we employed the approach [22] derived from hadronic scattering data, and adapted it for muonic atoms [23].
The SE correction has been improved compared to Ref. [11] and calculated within a fully rigorous QED framework, following Ref. [24]. Furthermore, we investigated the model dependence of SE, which resulted in increased uncertainties compared to the previous results. Similarly, for the relativistic recoil (RRecoil) correction, we adopted the approach of Ref. [25] and improved it by incorporating the 2pF model instead of the previously used exponential model.
Finally, the most cumbersome and challenging correction, the nuclear-polarization (NP) effect, has been calculated in the same manner as described in Ref. [13], now for all states appearing in the spectra. The uncertainties have been conservatively estimated as the spread of the results for the nine different Skyrme parametrizations that have been used in Ref. [13] and shown to cover a wide range in the parameter space. Special care had to be taken in the case of the $3d_{3/2}$ and $3d_{5/2}$ levels due to the resonances between the muonic transitions $3d_{3/2}\rightarrow 2p_{3/2}$, $3d_{5/2}\rightarrow 2p_{3/2}$ and $3d_{5/2}\rightarrow 2p_{1/2}$ and the lowest octupole nuclear excitation [26], see Sup. Mat. for further information.
In our calculations, we also evaluated the screening effect of the interaction between the muon and the outer surrounding electrons following Ref. [27], assuming the lowest electronic shell is filled as $(1s)^{2}$. Despite its relatively large absolute impact on individual energy levels, the screening effect remains constant across the low part of the muonic spectrum. Therefore, the number of remaining electrons and the exact electronic configuration do not affect the spectral lines of heavy muonic systems relevant for the rms radius evaluation.
In Table 1, we present all corrections, for simplicity calculated perturbatively (see Sup. Mat. for details) on the wavefunctions with FNS from 2pF, to illustrate the hierarchy of the effects. For the fitting procedure, all VP effects have been included in the Hamiltonian and, therefore, treated non-perturbatively.
Experimental data. — For our current evaluation, we have used the experimental measurements of $\mu-^{208}$b reported in Ref. [11] (see Table 2) and recently confirmed them to be reliable [28]. The data are split in three runs, which correspond to the high-, medium-, and low-energy parts of the spectrum. This split was arranged to allow for specific tuning of the detector. We performed the fit with respect to all transitions, but also analyzed the results from individual runs.
Set Transition Energy (keV)
Run I $2p_{3/2}-1s_{1/2}$ 5962.854(90)
$2p_{1/2}-1s_{1/2}$ 5778.058(100)
Run II $3d_{3/2}-2p_{1/2}$ 2642.332(30)
$3d_{5/2}-2p_{3/2}$ 2500.590(30)
$3d_{3/2}-2p_{3/2}$ 2457.569(70)
Run III $3p_{3/2}-2s_{1/2}$ 1507.754(50)
$3p_{1/2}-2s_{1/2}$ 1460.558(32)
$2s_{1/2}-2p_{1/2}$ 1215.330(30)
$2s_{1/2}-2p_{3/2}$ 1030.543(27)
Table 2: Experimentally measured transitions with their error bars [11].
Finally, it is important to note that the most recent spectral measurement of muonic lead was performed almost 40 years ago, and there continues to exist tension between theory and the experiment. With our rigorous QED calculation we aim to prompt a new experimental campaign to measure the spectrum of muonic lead with a focus on state-of-the-art uncertainty analysis explicitly splitting statistical and systematical errors.
Fitting procedure. — We calculated the transition energies as a function of 2pF parameters $a$ and $c$, including all corrections described above. The best-fit parameters are obtained by minimizing $\chi^{2}$ between the experimental data and the theory prediction. The theoretical uncertainties arise from a combination of model dependencies, deviations between different numerical methods, and empirical estimations of unaccounted higher-order contributions, and are fully taken into account. Experimental uncertainties are assumed to be uncorrelated as shown in Table 2. For more detailed information, see Sup. Mat.
Results and Discussion. — The best-fit 2pF parameters $c$ and $a$, together with the corresponding rms charge radius, are shown in Table 3. Fig. 2 depicts the likelihood contours for the 2pF parameters up to the $5\sigma$ level. The fit was performed for each of the 3 runs separately and the final value was extracted from a combined fit to all the experimental data. The individual fits are represented by different colours: purple, green, and blue correspond to runs I, II, and III, respectively. The full fit is depicted in orange. The insets show the residual transition-energy differences between best-fit theory and measurement for each of the fits. The ellipses corresponding to the 3 individual runs are in mild tension with each other which is ultimately responsible for the poor fit quality of the full fit ($\chi^{2}/\mathrm{DoF}=9.7$). Interestingly, the best-fit values for the individual runs still fall closely to a line of constant rms charge radius. Note that the figure does not show the best-fit value for Run I as it is far off the other two, but can be found in Table 3.
The likelihood contours for all the fits are well aligned with the lines of constant radius, which explains the relatively small uncertainty on the rms charge radius compared to the individual 2pF parameters. Interestingly, Run I, shown in purple, has a narrow but very extended likelihood surface, which results in a comparatively larger error of the rms charge radius than for the lower-energy runs. This is opposite to the usual notion that the high-energy transitions are most sensitive to the rms charge radius and can be explained by the large theory uncertainty for these low-lying states. We therefore believe that all experimental data should be utilized for the fitting procedure.
To understand the implications of the present analysis it is important to remember the three basic assumptions which underlie the present work:
  • (i)
    The experimental values including their uncertainties are correct.
  • (ii)
    We have correctly accounted for every relevant QED effect and the theory uncertainties.
  • (iii)
    The nuclear charge distribution is close enough to a 2pF model.
We lack the expertise and raw data to comment on the validity of (i). We have assessed assumption (ii) to the best of our knowledge. It is well known that muonic spectroscopy is very sensitive to nuclear properties, and thus we see inaccuracy of (iii) as the most likely explanation for the remaining discrepancy between individual runs.
Figure 2: The fit of the spectrum of $\mu-^{208}$b to a 2pF distribution. The colored regions show the $5\sigma$ likelihood for the 2pF parameters and the dotted contours highlight the $5$, $3$, and $1\sigma$ steps. Each colour corresponds to a fit of different groups of transitions: purple only fits Run I, green corresponds to Run II, blue to Run III, and orange is the combined fit of all the transitions. The insets show the resulting residual transition energies in the same colours, with measured transitions in darker shade and resulting fine-structure splittings in lighter colour. The previous fit as taken from Ref. [11] is shown in black wherever available, and the error bars include experimental errors only as the theory uncertainty was not explicitly quoted. The dashed lines correspond to curves of constant radius. The coloured circles indicate the best-fit parameters and the red circle the final values for the overall fit. Note that the best-fit value for Run I lies outside of the plotted parameter range, as indicated by the red arrow.
Comparison with previous data. — Comparison of our results with the most recent full extraction of the rms charge radius of lead in [11] is straightforward. The underlying data is the same; however, we have updated QED theory and error analysis. We find a best-fit rms charge radius which is in mild tension with their previously reported muonic spectroscopy value, see Table 3.
The results of Bergem [11] were combined with electron scattering data and optical isotope-shift measurements in the compilations of Fricke [3] and Angeli [16]. This combination of data sets is performed to empirically correct for the assumption of a nuclear 2pF distribution. Taken at face value, their final number is in strong tension with our result. We discuss below that the systematic error introduced through the exact shape of the nuclear charge density can not be quantified exactly; however we still believe that the tabulated rms charge radius of lead requires updating.
Model dependence. — Based on the recent indication that 208Pb is slightly deformed [29], we investigated the effect on the rms radius value. We calculated the corresponding deformation parameter $\beta$, following the methods described in Sup. Mat., and performed the fit for the deformed Fermi model. The resulting rms radii of the two models are the same, up to all significant figures reported in Table 3. The difference in $\chi^{2}/\mathrm{DoF}$ between the two models is also negligible.
As a basic test of model dependence we also considered a 3pF model with additional free parameter $w$ (see Sup. Mat.). The resulting fit is consistent with the 2pF model and the reduced $\chi^{2}/\mathrm{DoF}$ does not suggest preference for this modification. We would like to stress that a full analysis of the model dependence would require a parametrization of all possible density functions and is therefore not feasible until new, ab initio nuclear densities are available.
In the absence of ‘better’ test-distributions, rooted in physical principles, the exact level of accuracy at which the deviation of the nuclear charge distribution from 2pF affects the rms charge radius can only be estimated. A conservative limit would be to compare the 2pF specific value from muonic spectroscopy with values extracted from electron-scattering experiments. The latter are, in principle, sensitive to the shape of the charge distribution; however, the subtraction of QED effects faces challenges. As we lack the expertise to comment on the procedure, we simply highlight that the rms charge radii tend to deviate on the level of $\sim$0.001 fm.
Run I Run II Run III Total Ref. [11]
$c$ 6.55(17) 6.685(66) 6.639(7) 6.6409(25) 6.6447(5)
$a$ 0.59(11) 0.505(34) 0.531(5) 0.5283(18) 0.5249(3)
rms 5.525(46) 5.508(5) 5.509(2) 5.5062(5) 5.5040(11)
radius
/${}^{\chi^{2}}/_{\mathrm{DoF}}$   — 0.0001 2.9 9.5 187
Table 3: Nuclear parameters of ${}^{208}\mathrm{Pb}$ determined from the fits to muonic transition energies, in fm.
[…]
Contrary to conventional understanding, our analysis demonstrates that transitions involving $2s$, $2p$ and $3p$ states impose at least equally stringent constraints on radius extraction as traditionally emphasized $2p-1s$ transitions, which can be important for future experiments.
Our final value is consistent with the previous work in Ref. [11], which analyzed the same muonic spectral data, but lacks rigorous QED predictions for some of the effects. Compared to the accepted value, which combines muonic spectroscopy with electron-scattering measurements, we find a significantly larger rms charge radius. We are convinced that the current standard value might be too small and requires reconsideration in light of this work. Our results strongly indicate the necessity of a systematic reevaluation of the rms charge radii across the periodic table.
Acknowledgements. — The Authors thank A. Knecht, B. Ohayon, V. A. Yerokhin, and V. A. Zaytsev for discussions. N. S. O thanks ISOQUANT for financial support. This article comprises parts of the PhD thesis work of Z. M. and Z. S. to be submitted to Heidelberg University.